Faculty salaries for a random sample of teachers in the public-school system of a certain town were
coded by dividing each salary by 1000. Find the MAD and standard deviation of these salaries if the
coded observations are 18, 15, 21, 19, 13, 15, 14, 23, 18, and 16 pesos.
1
Expert's answer
2020-11-08T18:52:27-0500

Solution



Mean

"\bar{x}= {\sum x \over n} = {172000 \over 10} = 17200"

Mean Absolute Deviation

"MAD={1 \over n}\sum|x-\bar{x}|""={1 \over 10}(26000)""=2600"

Standard Deviation

"sd=\sqrt {\sum {(x- \bar{x})}^2 \over n-1}"

"=\sqrt{916000000 \over 9}""= 3190.2630"